the length of a rectangle is five times its width. If the perimeter of the rectangle is 60m find its area?
Answer by checkley79(3341)

P=2L+2W
L=5W
60=2(5W)+2W
60=10W+2W
60=12W
W=60/12
W=5M. ANS. FOR THE WIDTH.
L=5*5=25M. ANS. FOR THE LENGTH.
PROOF:
60=2*25+2*5
60=50+10
60=60
A=5*25
A=125M^2 ANS. FOR THE AREA.